OpenStudy (dls):

Hello,I'm working on collecting different and important maths series and expansions,please add on something if you know/like. PS:I'm preparing these for limits so accordingly I need some important ones that I may have left.

OpenStudy (dls):

I have expansions of most of the trigonometrical functions(sinx,cosx,tanx) as well as inverse trigonometrical functions(arcsin,arccos,arctan).I have =>e^x and e^(-x) =>log(1+x) and log(1-x) =>a^x =>(1+x)^(1/x)

OpenStudy (dls):

@Callisto

OpenStudy (anonymous):

don't forget geometric \(\frac{1}{1-x}\) and all its relatives, like \(\frac{1}{1+x}\), \(\frac{1}{x+2x^2}\) etc

OpenStudy (callisto):

Maybe \(\frac{1}{(1-x)^2}\) and \(\frac{2x}{(1-x)^3}\)? You can "derive" it from \(\frac{1}{1-x}\) though.

OpenStudy (dls):

that is all?

OpenStudy (callisto):

I don't know, since I rarely use series in finding the limits :(

OpenStudy (dls):

and i mostly use them :P anyway thanks!